# Once Welcomed, Then Scapegoated: The Enduring Consequences of Assimilation Policies in the Wake of Mass Migration

### v1

@@
PhD in PPA - Bocconi University, vinicius.dinizschuabb@phd.unibocconi.it
(September 2, 2025)
This paper examines the short- and long-term effects of immigrant assimilation policies in Brazil following the Mass Migration period. I focus on the Nationalization Campaign, launched by the federal government amid rising anti-immigrant sentiment during the Great Depression and the two World Wars. Using newly assembled archival data spanning the twentieth century, I assess the impact of these policies on the educational outcomes of immigrants and their descendants, following the forced closure of hundreds of immigrant-community schools. In the short term, the campaign significantly reduced educational attainment among targeted immigrant groups across state, municipal, and individual levels. In the long term, immigrants who were school-age during the campaign attained less education over their lifetimes, with adverse effects extending into the second generation. The magnitude of these effects was mediated by the degree of cultural proximity between immigrant groups and native Brazilians. These findings underscore the enduring consequences of assimilationist policies and offer insights for contemporary debates on immigration and education policy.
@@ 1 Introduction
A key mechanism driving these effects was cultural proximity, as highlighted by Fouka et al. (2021), Tabellini (2019), and Lecce et al. (2022). The Nationalization Campaign promoted a narrow vision of nationalism centered on the Portuguese language and Catholicism. Empirical evidence supports historical accounts indicating that immigrants from culturally distant backgrounds faced greater barriers to adaptation and, in some cases, actively resisted these norms. Japanese and German communities, for example, were more severely affected than Italian ones, particularly in cultural and educational domains. Cultural change tends to be gradual and often meets resistance, especially when majority and minority cultures diverge significantly (Giuliano and Nunn, 2021; Carvalho and Koyama, 2016). In this context, cultural and genetic proximity may condition the effectiveness of assimilation policies and help explain variation in immigrant integration outcomes (Arbatli et al., 2020; Spolaore and Wacziarg, 2009).
@@
The foundations of this process lie in concerns with demographic composition and cultural cohesion, ideas central to political thought since Plato and Aristotle, who viewed education as essential to the formation of the polis. [^1] See, for example, Plato’s Republic, Laws, and Statesman. Modern nation-building processes have similarly relied on assimilation policies, especially educational reforms, as tools to foster institutional trust and national identity (Alesina et al., 2021; Alesina and La Ferrara, 2002; Hall, 1992; Hobsbawm, 1992; John R. Lott, 1999). Authoritarian regimes, in particular, have strong incentives to pursue such policies,222“… the diminishing returns to controlling the citizenry through force increases the marginal return to indoctrination.” (John R. Lott, 1999) often adopting measures aimed at fostering trust between minority groups, the broader population, and national institutions (Alesina and La Ferrara, 2002).
@@
Brazil’s Nationalization Campaign exemplifies how state-led assimilation policies can shape long-term outcomes. This context provides a valuable setting to examine the intergenerational effects of forced assimilation on human capital formation across diverse immigrant groups (Becker and Tomes, 1986; Cunha and Heckman, 2007). Since intergenerational cultural transmission plays a key role in the accumulation of skills and knowledge (Bisin and Verdier, 2000; Bisin et al., 2021), assimilation policies may have led to persistent gaps in educational attainment (Botticini and Eckstein, 2012; Squicciarini, 2020; Becker and Woessmann, 2009). These disparities may, in turn, affect long-run local development, as immigrant-linked human capital has been shown to yield lasting economic benefits (Rocha et al., 2017; Becker et al., 2011).
@@
This analysis contributes to our understanding of the mechanisms driving intergenerational inequality and local development under state-led immigration and education policies (Abramitzky et al., 2021; Torche and Ribeiro, 2010; Bloome et al., 2018). It also adds to the growing literature on assimilation efforts implemented through educational reforms, particularly those involving the imposition of a national language of instruction (Fouka, 2020; Lleras-Muney and Shertzer, 2015; Clots-Figueras and Masella, 2013; Eriksson, 2014).
@@ 2 Historical Context
The Mass Migration of the nineteenth and twentieth centuries brought a diverse array of ethnic groups to the Americas (Acosta Arcarazo, 2018; Sánchez-Alonso, 2019; Fitzgerald and Cook-Martín, 2014), with Brazil receiving nearly 5 million European and Asian immigrants between 1819 and 1947, one of the largest migration episodes in modern history (Balderas and Greenwood, 2010). Many of these immigrants settled in the economically developing southern regions of Brazil, particularly in remote rural areas of states like Rio Grande do Sul and São Paulo, where they established largely autonomous communities. Within these communities, immigrant-run schools were central institutions, serving to transmit language, religion, and cultural traditions across generations, especially in areas underserved by the national education system (Kreutz, 2010). For instance, by 1940, Brazil counted 644,255 German-speaking and 458,093 Italian-speaking residents.
@@ 2.1 Migration Policies
The Great Depression of 1929 triggered a severe economic crisis in Brazil, where coffee exports—central to the national economy—lost 70% of their value. In the wake of this collapse, political leaders such as Getúlio Vargas began to scapegoat immigrants, accusing them of taking jobs from native Brazilians. Immigrants, once welcomed to replace enslaved labor, were now portrayed as contributing to national instability. Rising xenophobia during the 1930s and 1940s coincided with growing public support for restrictive measures, reshaping the state’s approach to immigration.
@@ 4.2.1 Empirical Strategy: Interaction with Information on Rural Schools
$\Delta\;Education_{m,1950-1940}=$ $\alpha+\beta\;(Non{\text{-}}Southern\;European\;Immi._{m,1940}\;*\;\text{Rural Schools}_{m,1940})\;+$
$+\delta\;Non{\text{-}}Southern\;European\;Immi._{m,1940}+\rho\;Rural\;Schools_{m,1940}\;+$
@@ 5 Long-term and intergenerational effects
Consistent with evidence from other contexts (Carvalho and Koyama, 2016; Fouka, 2020; Lleras-Muney and Shertzer, 2015), the Brazilian Nationalization Campaign had significant adverse effects on school-age immigrant children and their families. Given the strong link between intergenerational and cross-family inequality (Becker and Tomes, 1986), such disruptions are likely to generate lasting consequences. To assess these long-term effects, I use individual-level data from the 1991, 2000, and 2010 Brazilian Censuses. Beginning in 1991, the Census recorded both immigrants’ countries of origin and their year of arrival in Brazil—key variables for the empirical strategy. These household datasets also allow the identification of second-generation immigrants, defined as individuals living with at least one foreign-born parent, enabling an intergenerational analysis of educational outcomes. The results reveal persistent gaps in attainment between immigrants, and their descendants, exposed to the Nationalization Campaign (1938–1945) and comparable cohorts of immigrant origin who completed their schooling in their countries of origin.
@@ 6 Discussion and Conclusion
The findings are consistent with evidence from other historical contexts (Carvalho and Koyama, 2016; Fouka, 2020; Lleras-Muney and Shertzer, 2015) and underscore the cumulative and intergenerational impacts of early-life policy shocks. Cultural proximity played a central role: immigrants from Southern Europe, such as Italians, faced relatively smaller educational setbacks, likely due to linguistic and cultural similarities with the dominant Brazilian population. By contrast, Japanese and German immigrants, who faced greater cultural distance, suffered substantial and lasting educational penalties.

### v2

@@
PhD in Public Policy and Administration, June 2025, Bocconi University, Italy (vinicius.dinizschuabb@phd.unibocconi.it); Research Fellow at the Institute for Mobility and Social Development (IMDS) and Research Associate/Postdoctoral researcher at the Center for Empirical Studies in Economics (FGV CEEE), Rio de Janeiro, Brazil.I am grateful to Maristella Botticini (Bocconi University) for her supervision and guidance. I thank Laura Ogliari and Giampaolo Lecce (University of Bergamo) for helpful comments, and Vincenzo Galasso, Paolo Pinotti, Manoel Gehrke, Henrique Alpalhão, Paolo Agnolin, Elena Marseglia, and Gian Maria Mallarino (Bocconi University) for valuable suggestions. I also thank Claudio Ferraz (UBC and PUC-Rio), Pedro Américo (USP), and Leonardo Monasterio (IPEA) for insightful feedback on earlier stages of this project. Finally, I thank the Graeff family for the inspiration provided during my exchange semester in Karlsruhe, Germany, in 2012.
(For the most up-to-date version, see: arxiv.org/abs/2509.02836)
This paper examines the short- and long-term effects of immigrant assimilation policies in Brazil after the Mass Migration period. It focuses on the Nationalization Campaign, launched amid rising anti-immigrant sentiment during the Great Depression and the World Wars. Using newly assembled archival data spanning the twentieth century, I assess the campaign’s impact on education among immigrants and their descendants following the closure of immigrant-community schools. In the short term, the policy sharply reduced educational attainment among targeted groups; in the long term, school-age immigrants during the campaign completed less education over their lifetimes, with adverse effects persisting into the second generation. The magnitude of these impacts varied with cultural proximity to native Brazilians.
@@ 1 Introduction
A key mechanism driving these effects was cultural proximity, as highlighted by Lecce et al. (2022), Fouka et al. (2021), and Tabellini (2019). The Nationalization Campaign promoted a narrow vision of nationalism centered on the Portuguese language and Catholicism. Empirical evidence supports historical accounts indicating that immigrants from culturally distant backgrounds faced greater barriers to adaptation and, in some cases, actively resisted these norms. Japanese and German communities, for example, were more severely affected than Italian ones, particularly in cultural and educational domains. Cultural change tends to be gradual and often meets resistance, especially when majority and minority cultures diverge significantly (Giuliano and Nunn, 2021; Carvalho and Koyama, 2016). In this context, cultural and genetic proximity may condition the effectiveness of assimilation policies and help explain variation in immigrant integration outcomes (Arbatli et al., 2020; Spolaore and Wacziarg, 2009).
@@
The foundations of this process lie in concerns with demographic composition and cultural cohesion, ideas central to political thought since Plato and Aristotle, who viewed education as essential to the formation of the polis. [^1] See, for example, Plato’s Republic, Laws, and Statesman. Modern nation-building processes have similarly relied on assimilation policies, especially educational reforms, as tools to foster institutional trust and national identity (Alesina et al., 2021; Alesina and La Ferrara, 2002; John R. Lott, 1999; Hall, 1992; Hobsbawm, 1992). Authoritarian regimes, in particular, have strong incentives to pursue such policies,222“… the diminishing returns to controlling the citizenry through force increases the marginal return to indoctrination.” (John R. Lott, 1999) often adopting measures aimed at fostering trust between minority groups, the broader population, and national institutions (Alesina and La Ferrara, 2002).
@@
Brazil’s Nationalization Campaign exemplifies how state-led assimilation policies can shape long-term outcomes. This context provides a valuable setting to examine the intergenerational effects of forced assimilation on human capital formation across diverse immigrant groups (Cunha and Heckman, 2007; Becker and Tomes, 1986). Since intergenerational cultural transmission plays a key role in the accumulation of skills and knowledge (Bisin et al., 2021; Bisin and Verdier, 2000), assimilation policies may have led to persistent gaps in educational attainment (Squicciarini, 2020; Botticini and Eckstein, 2012; Becker and Woessmann, 2009). These disparities may, in turn, affect long-run local development, as immigrant-linked human capital has been shown to yield lasting economic benefits (Rocha et al., 2017; Becker et al., 2011).
@@
This analysis contributes to our understanding of the mechanisms driving intergenerational inequality and local development under state-led immigration and education policies (Abramitzky et al., 2021; Bloome et al., 2018; Torche and Ribeiro, 2010). It also adds to the growing literature on assimilation efforts implemented through educational reforms, particularly those involving the imposition of a national language of instruction (Fouka, 2020; Lleras-Muney and Shertzer, 2015; Eriksson, 2014; Clots-Figueras and Masella, 2013).
@@ 2 Historical Context
The Mass Migration of the nineteenth and twentieth centuries brought a diverse array of ethnic groups to the Americas (Sánchez-Alonso, 2019; Acosta Arcarazo, 2018; Fitzgerald and Cook-Martín, 2014), with Brazil receiving nearly 5 million European and Asian immigrants between 1819 and 1947, one of the largest migration episodes in modern history (Balderas and Greenwood, 2010). Many of these immigrants settled in the economically developing southern regions of Brazil, particularly in remote rural areas of states like Rio Grande do Sul and São Paulo, where they established largely autonomous communities. Within these communities, immigrant-run schools were central institutions, serving to transmit language, religion, and cultural traditions across generations, especially in areas underserved by the national education system (Kreutz, 2010). For instance, by 1940, Brazil counted 644,255 German-speaking and 458,093 Italian-speaking residents.
@@ 2.1 Migration Policies
The Great Depression of 1929 triggered a severe economic crisis in Brazil, where coffee exports—central to the national economy, lost 70% of their value. In the wake of this collapse, political leaders such as Getúlio Vargas began to scapegoat immigrants, accusing them of taking jobs from native Brazilians. Immigrants, once welcomed to replace enslaved labor, were now portrayed as contributing to national instability. Rising xenophobia during the 1930s and 1940s coincided with growing public support for restrictive measures, reshaping the state’s approach to immigration.
@@ 4.2.1 Empirical Strategy: Interaction with Information on Rural Schools
$\Delta\;Educ._{m,1950-1940}=$ $\alpha+\beta\;(Non{\text{-}}Southern\;European\;Imm._{m,1940}\;*\;\text{Rural Schools}_{m,1940})\;+$
$+\delta\;Non{\text{-}}Southern\;European\;Imm._{m,1940}+\rho\;Rural\;Schools_{m,1940}\;+$
@@ 5 Long-term and intergenerational effects
Consistent with evidence from other contexts (Fouka, 2020; Carvalho and Koyama, 2016; Lleras-Muney and Shertzer, 2015), the Brazilian Nationalization Campaign had significant adverse effects on school-age immigrant children and their families. Given the strong link between intergenerational and cross-family inequality (Becker and Tomes, 1986), such disruptions are likely to generate lasting consequences. To assess these long-term effects, I use individual-level data from the 1991, 2000, and 2010 Brazilian Censuses. Beginning in 1991, the Census recorded both immigrants’ countries of origin and their year of arrival in Brazil—key variables for the empirical strategy. These household datasets also allow the identification of second-generation immigrants, defined as individuals living with at least one foreign-born parent, enabling an intergenerational analysis of educational outcomes. The results reveal persistent gaps in attainment between immigrants, and their descendants, exposed to the Nationalization Campaign (1938–1945) and comparable cohorts of immigrant origin who completed their schooling in their countries of origin.
@@ 6 Concluding Remarks
The findings are consistent with evidence from other historical contexts (Fouka, 2020; Carvalho and Koyama, 2016; Lleras-Muney and Shertzer, 2015) and underscore the cumulative and intergenerational impacts of early-life policy shocks. Cultural proximity played a central role: immigrants from Southern Europe, such as Italians, faced relatively smaller educational setbacks, likely due to linguistic and cultural similarities with the dominant Brazilian population. By contrast, Japanese and German immigrants, who faced greater cultural distance, suffered substantial and lasting educational penalties.